OpenStudy (anonymous):

how do i solve this question? 4x+4y=16 2x-2y=16 systems of equations are very confusing

OpenStudy (anonymous):

it says i have to find the solution by using elimination by addition

OpenStudy (amistre64):

Solve for one variable and stick it into the equation. Let go with and solve for (y): 4x+4y=16 4y=-4x+16 y = -x + 4 now that we know what y equals, plug it into the other equation. 2x -2(-x+4)=16 2x +2x -8= 16 4x = 24 x= 6 We now have a solid value for "x" and can put into the equation that we found for "y" y = -x +4 y = -(6) +4 y= -2 (6,-2) is the solution to the system of equations.

OpenStudy (anonymous):

ok thank you

OpenStudy (amistre64):

There are a few ways to find it, but it is best to stick with whatever method you know best and are comfortable with. A test just wants the right answer, not how you get it :)

OpenStudy (anonymous):

If you want to solve by addition then: You have to add a multiple of the second equation to the first equation to try to cancel out the variables. Then carry on with regular algebra to solve for the other variable.

OpenStudy (anonymous):

Edit -> cancel out ONE of the variables...